# Bulk-Boundary Correspondence and Exceptional Points for a Dimerized Hatano-Nelson Model with Staggered Potentials

### v1

@@ III.3 Winding numbers
The eigenvalues of a non-Hermitian Hamiltonian are in general complex, and can trace a closed path around an arbitrary reference complex energy $E_{B}$, in which case a winding number can be assigned [36, 37, 18, 38, 32],
[…]
provided that $H(k)$ has a point gap, that is, $\det(H(k)-E_{B})\neq 0$. For concreteness we will chose $E_{B}=0$. Observe that the definition (III.30) is universal since no symmetry requirements are made. Therefore, one dimensional non-Hermitian systems—in contrast to Hermitian ones—can have non-trivial topology in the absence of any spatial or non-spatial symmetries.
If symmetries are present, then they lead to additional topological invariants [14]. So far, mostly the case of a two-band model with sub-lattice symmetry has been considered [39, 40, 41, 42, 11, 38]. The Hamiltonian of such a model takes the form
$H(k)=\left(\begin{array}[]{cc}0&\hat{H}_{1}(k)\\ \hat{H}_{2}(k)&0\\ \end{array}\right)\,.$ (III.31)
@@
[…]
@@
$I_{1}=\nu_{1}+\nu_{2},$ (III.33)
which is consistent with the definition of $I_{1}$ in Eq. (III.30), and
$I_{2}=\nu_{2}-\nu_{1}$ (III.34)
as independent winding numbers. For our model with $u=0$, we have $H_{j}(k)=|H_{j}(k)|e^{\text{i}\phi_{j}}$ and evaluating the integrals (III.33,III.34) leads to
$I_{1}(u=0)=\begin{cases}-1,&\text{if}\quad~{}\left|\frac{W_{L}}{V_{R}}\right|>1\land\left|\frac{V_{L}}{W_{R}}\right|>1\\ 1,&\text{if}\quad\left|\frac{W_{L}}{V_{R}}\right|<1\land\left|\frac{V_{L}}{W_{R}}\right|<1\\ 0,&\text{otherwise}\end{cases}$ (III.35)
@@
$I_{2}=\begin{cases}-2&\text{if}\quad~{}\left|\frac{V_{L}}{W_{R}}\right|<1\land\left|\frac{W_{L}}{V_{R}}\right|>1\\ 0&\text{if}\quad\left|\frac{V_{L}}{W_{R}}\right|>1\land\left|\frac{W_{L}}{V_{R}}\right|<1\\ -1&\text{otherwise}.\end{cases}$ (III.36)
Clearly, the winding numbers are not defined when $\left|\frac{V_{L}}{W_{R}}\right|=\left|\frac{W_{L}}{V_{R}}\right|=1$ which are precisely the gapless/exceptional lines discussed in the previous subsection. The two winding numbers $\nu_{1,2}$ are shown in the left panel of Fig. 2 for parameters $V_{R}=W_{R}=1$ (no loss of generality) in each phase as a function of $V_{L}$ and $W_{L}$.
Next, we briefly discuss the $\mathcal{PT}$-symmetric case. The Hamiltonian (III.3) for $V_{L}=V_{R}=V$ and $W_{L}=W_{R}=W$ satisfies the $\mathcal{PT}$ symmetry $\sigma^{x}H(k)\sigma^{x}=H(k)^{*}$ as well as chiral symmetry $\sigma^{z}H(k)^{\dagger}\sigma^{z}=-H(k)$, and this case is part of the AZ symmetry class AIII [14]. The phase diagram for $u=1$ is shown in Fig. 2(c) and consists of gapped phases and phases where exceptional points become dense in the thermodynamic limit. These phases will be discussed further in Sec. VI when we analyze the entanglement spectrum.
@@ IV.1 Diagonalization
Let us also recall the diagonalization of $\mathcal{T}_{N}$ (II) for open boundary conditions (OBC), that is, $\gamma=0$. In this case, $\mathcal{T}_{N}$ is an asymmetric tridiagonal 2-Toeplitz matrix whose eigenvalue problem was solved from different perspectives in Refs. [44, 45]. In App. A, we provide some details on both approaches. The connection between these two approaches was pointed out in Ref. [46] and the method [45] was previously applied to the Hermitian SSH chain in Refs. [47, 48].
The solution of the open boundary case can also be obtained by a simple transformation to the standard (symmetric) SSH chain [12], namely,
$S\mathcal{T}_{N}S^{-1}=\mathcal{T}_{N}^{SSH}$ (IV.1)
@@
$s_{2j-1}=\frac{V_{R}^{\frac{j-1}{2}}W_{R}^{\frac{j-1}{2}}}{V_{L}^{\frac{j-1}{2}}W_{L}^{\frac{j-1}{2}}},\quad s_{2j}=\frac{V_{R}^{\frac{j}{2}}W_{R}^{\frac{j}{2}}}{V_{L}^{\frac{j}{2}}W_{L}^{\frac{j}{2}}}\frac{W_{L}}{W_{R}},$ (IV.2)
@@
$V_{\text{eff}}=\sqrt{V_{L}V_{R}},\quad W_{\text{eff}}=\sqrt{W_{L}W_{R}}.$ (IV.3)
@@
which gives $N-1$ eigenvalues from (IV.4) and (IV.8). Together with the isolated root $\epsilon=\text{i}u$, one has all the $N$ eigenvalues of the tridiagonal matrix when $\gamma=0$.
@@
We remark that the transcendental equation (IV.9) has the same form as the transcendental equation of the quantum Ising chain with transverse field $\lambda=\sqrt{\delta}$ and $L/2$ sites [49]. Meanwhile, the eigenvector components can be written as
@@ IV.2.1 Odd N
Therefore, for non-vanishing hopping parameters, there are $N-1$ exceptional points. For finite $N$, these points always have an imaginary part. In the thermodynamic limit, however, the exceptional points will converge to the real axis
@@
×$\quad\times\left(\sqrt{\delta}+\frac{1}{\sqrt{\delta}}(1-\bar{u}^{2})+2\cos(\theta_{m})\right)\,.$ (IV.19)
The minimal gap (IV.15) as a function of $\bar{u}$ and $\delta$ is shown in Fig. 3.
Figure 3: The rescaled minimal gap $\Delta$ as a function of $\sqrt{\delta}$ and $\bar{u}$ for $N=101$ and open boundary conditions.
The exceptional points are therefore now given by
@@
Interestingly, the expression (IV.20) implies that real exceptional points do exist for finite $N$ provided that $\bar{u}\geq|\sin(\theta_{m})|$ or $\bar{u}\leq-|\sin(\theta_{m})|$. In the thermodynamic limit, critical fields $\bar{u}\rightarrow\pm 1$ emerge when $\theta_{m}$ approaches 00 or $\pi$. In Fig. 4 some examples of the loci of exceptional points for different values of the rescaled field $\bar{u}$ are shown.
Figure 4: Loci of exceptional points $\left(\sqrt{\delta}\right)_{\text{EP}}$ for $N=41$ and different rescaled fields $\bar{u}$.
Recall that when $N$ is odd, there is, in addition, an isolated root $\epsilon=\text{i}u$. We can impose that this root coincides with the quasienergies (IV.4), that is, $\text{i}u=\epsilon^{\pm}$. For non-zero hopping parameters, this implies again the condition (IV.16), but now associated with the degenerate quasienergy $\text{i}u$.
@@ IV.2.2 Even N
but with $\theta$ now given by by (IV.9). Since Eq. (IV.9) also involves $\sqrt{\delta}$, one has to consider both equations simultaneously. Note that both conditions are satisfied if any of the hopping parameters vanishes similar to the case of $N$ odd. Otherwise, after some manipulation, the intersection of (IV.21) and (IV.9) leads to,
@@
The transcendental equation (IV.25) is clearly independent of the field $u$. It can be solved numerically; there are $N-2$ solutions [50] with $\text{Re}(\theta_{\text{EP}}^{\prime})\geq 0$ which fully describe the field independent part of the discriminant. The numerical solution for different lattice sizes is shown in the upper panel of Fig. 6. Note that in this case the associated repeated quasienergy is not zero. Instead, they trace interesting curves in the complex plane, see the examples shown in the bottom panel of Fig. 6.
@@ V Bulk-boundary correspondence
It is well-known that the standard bulk-boundary correspondence for topological Hermitian systems breaks down in the non-Hermitian case. Mathematically, the fundamental issue is that the eigenvalues of a finite non-Hermitian matrix do, in general, not converge to the eigenvalues of the semi-infinite case. This is similar to a discontinuous point. More specifically, in the Hermitian case we can connect the topological invariant calculated for a periodic system with the number of edge states in a finite system. The energy of these edge states converges exponentially to zero with system size. This is what is known as the standard bulk-boundary correspondence.
In the non-Hermitian case, the eigenspectrum cannot be used to construct a bulk boundary correspondence for a finite system. Instead, it is known from Toeplitz theory that the singular value spectrum has to be used. The K-splitting theorem states [51] that a system with winding number $I$ has $K\geq|I|$ singular values which will converge to zero with increasing system size and which belong to states which are exponentially localized at the boundary. In the semi-infinite case these states become exact eigenstates with exactly zero energy but for finite system size they only get mapped exponentially close to zero by the Hamiltonian but are not exact eigenstates. I.e., in a finite non-Hermitian system the zero energy edge modes are ’hidden’, in general, and constitute very long-lived states. The proper bulk-boundary correspondence for non-Hermitian systems was recently fully explored in Ref. [25]. We note that this formalism does encompass the Hermitian case because for a Hermitian matrix the singular values are just the absolute values of the eigenvalues. Here we want to demonstrate the bulk-boundary correspondence explicitly for the model under consideration.
@@ V.1 Singular values
with boundaries fixed by the last line in (V.1). Using the method reviewed in App. A, we find that there are two solutions to the difference equations (V.1).
@@
The spectrum of (V.1) is given by the union of $\sigma_{\phi}^{2}$ and $\sigma_{\omega}^{2}$. Remarkably, the pairs of variables $\{V_{L},W_{R}\}$ and $\{V_{R},W_{L}\}$ decouple from each other. This means that the characteristic polynomial associated with (V.1) factorizes into two independent factors depending solely on each of these pairs.
The transcendental equations (V.7) and (V.11) can be solved numerically. We find that (V.7) has $N/2$ real solutions in the interval $0<\phi<\pi$ when $|V_{L}/W_{R}|\geq 1$. When $|V_{L}/W_{R}|<1$ a complex conjugate pair emerges near $\pi$. This complex root produces an exponentially small singular value $\sigma_{\phi}$. Similarly, we find that (V.11) has $N/2$ real solutions in the interval $0<\omega<\pi$ when $|V_{R}/W_{L}|\geq 1$ and $N/2-1$ real solutions plus a complex conjugate pair solution when $|V_{R}/W_{L}|<1$. In Fig. 7, we plot the smallest two singular values coming from each transcendental equation considering a lattice with $N=100$ sites.
@@
The Bloch Hamiltonian for the corresponding periodic problem with $u=0$ has block structure, see Eq. (III.31). There are therefore two winding numbers $\nu_{1,2}$ defined in Eq. (III.32). The k-splitting theorem then predicts that there are $K\geq|\nu_{1}|+|\nu_{2}|$ singular values $\sigma$ with $\lim_{N\to\infty}\sigma=0$ [51, 25]. This is fully consistent with our results, compare Fig. 7 with the phase diagram in the left panel of Fig. 2.
@@ V.2 Topological protection and hidden zero modes
In the limit of a semi-infinite chain, the system will have $K\geq|\nu_{1}|+|\nu_{2}|$ topologically protected edge modes with exactly zero energy. However, these states will only be eigenstates in this limit but not for a finite system. For a finite system, the singular values have to be considered as we have done above. Here we want to show explicitly how the exact edge states in the semi-infinite limit arise and that they become ’hidden’ if the system is finite.
@@
The condition $Hv=0$ leads to $v_{2}=0$ and $W_{R}v_{2j}+V_{L}v_{2j+2}=0$ implying that all even vector components are zero, $v_{2j}=0$. For the odd vector components we find the recurrence relation $V_{R}v_{2j-1}+W_{L}v_{2j+1}=0$ with the solution $v_{2j+1}=\left(-\frac{V_{R}}{W_{L}}\right)^{j}v_{1}$. In addition, we have to demand that the solution is normalizable, $||v||^{2}=|v_{1}|^{2}\sum_{j=0}^{\infty}\left(-\frac{V_{R}}{W_{L}}\right)^{2j}=1$. This is possible if $|W_{L}|>|V_{R}|$. In this case, the semi-infinite chain has a zero energy edge mode $v$ exponentially localized at the left edge with vector components
@@
Again, for a finite system the only solution to $\tilde{H}w=0$ is $w=0$. The non-trivial edge mode with energy zero only exists in the thermodynamic limit. We note that we find the exact same separation of parameters as for the singular values: the first zero mode depends only on $\{V_{R},W_{L}\}$ and the second zero mode only on $\{V_{L},W_{R}\}$. For a finite system, we can truncate the vectors $v,w$. They are then no longer eigenvectors, $Hv\neq 0$, but they do get mapped exponentially close to zero, i.e. $||Hv||\to 0$ for $N\to\infty$ and similarly for $w$. This is what we mean by hidden zero modes.
For the example considered in Fig. 8 with $V_{R}=W_{R}=1$ this means that the first zero mode is stable for $-1<V_{L}<1$ and the second zero mode for $|W_{L}|>1$. This is fully consistent with the singular values as well as with the winding numbers as is expected based on the K-splitting theorem. We thus do have explicitly confirmed the bulk-boundary correspondence in this case.
To check that the singular values with $\lim_{N\to\infty}\sigma=0$ are topologically protected while zero eigenvalues are, in general, not protected for a finite non-Hermitian system, we also show in Fig. 8 both quantities when calculated for a finite Hamiltonian $H$ ($N=200$) with a random complex matrix $A$ with elements $|a_{ij}|\leq 0.02$ added. The results show that the singular values are stable and thus indeed topologically protected while the eigenvalues are, in general, not protected. More specifically, we find that for the case $W_{L}=0.5$ shown in the lower left panel there are two degenerate zero eigenvalues in the unperturbed system for $-0.4\lesssim V_{L}\lesssim 0.5$. However, the small perturbation immediately moves them away from zero energy. They are not protected.
The situation is slightly more complicated in the case $W_{L}=1.5$ shown in the lower right panel of Fig. 8. In the unperturbed case, there are two degenerate zero energy eigenvalues for $-1.3\lesssim V_{L}\lesssim 1.5$. Interestingly, adding a perturbation does not completely remove them from zero energy. Instead, they are stable for $-1<V_{L}<1$. This can be understood as follows: in this regime, the non-Hermitian Hamiltonian with $V_{R}=W_{R}=1$ and $W_{L}=1.5$ can be adiabatically connected to a chiral Hermitian Hamiltonian without closing the gap. This can be achieved, for example, by $W_{R}\to 1.5$ and then $V_{R}\to V_{L}$. For the obtained Hermitian Hamiltonian one can calculate the winding number of the upper block which turns out to be $I=1$. The standard bulk-boundary correspondence then predicts the existence of two protected zero-energy edge modes. Since the gap never closes and the sub-lattice symmetry remains intact, these zero modes survive in the non-Hermitian case. However, the full topology of the non-Hermitian model also in this case is only captured by the singular values, not by the eigenvalues.

### v3

@@ III.3 Winding numbers
To characterize the different phases we have found, we can use topological invariants. The eigenvalues of a non-Hermitian Hamiltonian are in general complex and can trace, for periodic boundary conditions, a closed path around an arbitrary reference complex energy $E_{B}$, in which case a winding number can be assigned [41, 42, 20, 43, 37],
$I_{1}=\frac{1}{2\pi\text{i}}\int_{0}^{2\pi}dk~\log\det(H(k)-E_{B}),$ (III.29)
provided that $H(k)$ has a point gap at $E_{B}$, that is, $\det(H(k)-E_{B})\neq 0$. For concreteness we will chose $E_{B}=0$. Observe that the definition (III.29) is universal since no symmetry requirements are made. Therefore, one dimensional non-Hermitian systems—in contrast to Hermitian ones—can have non-trivial topology in the absence of any spatial or non-spatial symmetries.
If symmetries are present, then they lead to additional topological invariants [16, 22]. So far, mostly the case of a two-band model with sub-lattice symmetry has been considered [44, 45, 46, 47, 12, 43]. The Hamiltonian of such a model takes the form
$H(k)=\left(\begin{array}[]{cc}0&\hat{H}_{1}(k)\\ \hat{H}_{2}(k)&0\\ \end{array}\right)\,.$ (III.30)
@@
$\nu_{j}=\frac{1}{2\pi\text{i}}\int dk~{\rm tr}\,\left(q_{j}^{\dagger}\partial_{k}q_{j}\right)=\frac{1}{2\pi\text{i}}\int dk~\partial_{k}\log\det q_{j}\,.$ (III.31)
@@
$I_{1}=\nu_{1}+\nu_{2},$ (III.32)
which is consistent with the definition of $I_{1}$ in Eq. (III.29), and
$I_{2}=\nu_{2}-\nu_{1}$ (III.33)
as independent winding numbers. Note that in the Hermitian case, $\hat{H}_{2}(k)=\hat{H}_{1}^{\dagger}(k)$, we have $\nu_{2}=-\nu_{1}$ implying that $I_{1}\equiv 0$ and only one non-trivial winding number, $I_{2}$, remains. For our model with $u=0$, we can write $H_{j}(k)=|H_{j}(k)|e^{\text{i}\phi_{j}}$ and evaluating the integrals (III.32,III.33) leads to
$I_{1}(u=0)=\begin{cases}-1,&\text{if}\quad~\left|\frac{W_{L}}{V_{R}}\right|>1\land\left|\frac{V_{L}}{W_{R}}\right|>1\\ 1,&\text{if}\quad\left|\frac{W_{L}}{V_{R}}\right|<1\land\left|\frac{V_{L}}{W_{R}}\right|<1\\ 0,&\text{otherwise}\end{cases}$ (III.34)
@@
$I_{2}=\begin{cases}-2&\text{if}\quad~\left|\frac{V_{L}}{W_{R}}\right|<1\land\left|\frac{W_{L}}{V_{R}}\right|>1\\ 0&\text{if}\quad\left|\frac{V_{L}}{W_{R}}\right|>1\land\left|\frac{W_{L}}{V_{R}}\right|<1\\ -1&\text{otherwise}.\end{cases}$ (III.35)
Clearly, the winding numbers are not defined when $\left|\frac{V_{L}}{W_{R}}\right|=\left|\frac{W_{L}}{V_{R}}\right|=1$ which are precisely the gapless/exceptional lines discussed in the previous subsection. The two winding numbers $\nu_{1,2}=(I_{1}\mp I_{2})/2$ are shown in the left panel of Fig. 2 in each phase as a function of $V_{L}$ and $W_{L}$.
@@ IV.1 Diagonalization
Let us also recall the diagonalization of $\mathcal{T}_{N}$ (II) for open boundary conditions (OBC), that is, $\gamma=0$. In this case, $\mathcal{T}_{N}$ is an asymmetric tridiagonal 2-Toeplitz matrix whose eigenvalue problem was solved from different perspectives in Refs. [49, 50]. In App. B, we provide some details on both approaches. The connection between these two approaches was pointed out in Ref. [51] and the method [50] was previously applied to the Hermitian SSH chain in Refs. [52, 53].
The solution of the open boundary case can also be obtained by a simple similarity transformation to the standard (symmetric) SSH chain which leaves the spectrum invariant [14], namely,
$S^{-1}\mathcal{T}_{N}S=\mathcal{T}_{N}^{SSH}$ (IV.1)
@@
$s_{2j-1}=\frac{V_{R}^{\frac{j-1}{2}}W_{R}^{\frac{j-1}{2}}}{V_{L}^{\frac{j-1}{2}}W_{L}^{\frac{j-1}{2}}},\quad s_{2j}=\frac{V_{R}^{\frac{j}{2}}W_{R}^{\frac{j}{2}}}{V_{L}^{\frac{j}{2}}W_{L}^{\frac{j}{2}}}\frac{W_{L}^{\frac{1}{2}}}{W_{R}^{\frac{1}{2}}},$ (IV.2)
@@
$V_{\text{eff}}=\sqrt{V_{L}}\sqrt{V_{R}},\quad W_{\text{eff}}=\sqrt{W_{L}}\sqrt{W_{R}}.$ (IV.3)
@@
which gives N1N-1 eigenvalues from (IV.4) and (IV.8). Together with the isolated root $\epsilon=\text{i}u$, one has all the $N$ eigenvalues of the tridiagonal matrix when $\gamma=0$.
@@
We remark that the transcendental equation (IV.9) has the same form as the transcendental equation of the quantum Ising chain with transverse field $\lambda=\sqrt{\delta}$ and $N/2$ sites [54]. Meanwhile, the eigenvector components can be written as
@@ IV.2.1 Odd N
Therefore, for non-vanishing hopping parameters, there are N1N-1 exceptional points associated with the eigenvalue $\epsilon=0$. For finite $N$, these points always have an imaginary part. In the thermodynamic limit, however, the exceptional points will converge to the real axis
@@
$\quad\times\left(\sqrt{\delta}+\frac{1}{\sqrt{\delta}}(1-\bar{u}^{2})+2\cos(\theta_{m})\right)\,.$ (IV.19)
The exceptional points, which are still equivalent to the points where the gap closes, are therefore now given by
@@
Interestingly, the expression (IV.20) implies that now real exceptional points do exist even for finite $N$ provided that $\bar{u}\geq|\sin(\theta_{m})|$ or $\bar{u}\leq-|\sin(\theta_{m})|$. In the thermodynamic limit, critical fields $\bar{u}\rightarrow\pm 1$ emerge when $\theta_{m}$ approaches 0 or $\pi$. In Fig. 4 some examples of the loci of exceptional points for different values of the rescaled field $\bar{u}$ are shown.
Figure 4: Loci of exceptional points $\left(\sqrt{\delta}\right)_{\text{EP}}$, see Eq. (IV.20), for $N=41$ and different rescaled fields $\bar{u}$.
Recall that when NN is odd, there is, in addition, an isolated root $\epsilon=\text{i}u$. We can impose that this root coincides with the quasienergies (IV.4), that is, $\text{i}u=\epsilon^{\pm}$. For non-zero hopping parameters, this implies again the condition (IV.16), but now associated with the degenerate quasienergy $\text{i}u$.
@@ IV.2.2 Even N
but with $\theta$ now given by (IV.9). Since Eq. (IV.9) also involves $\sqrt{\delta}$, one has to consider both equations simultaneously. Note that both conditions are satisfied if any of the hopping parameters vanishes similar to the case of $N$ odd. Otherwise, after some manipulation, the intersection of (IV.21) and (IV.9) leads to,
@@
The transcendental equation (IV.25) is clearly independent of the field $u$. It can be solved numerically; there are N2N-2 solutions [55] with $\text{Re}(\theta_{\text{EP}}^{\prime})\geq 0$ which fully describe the field independent part of the discriminant. The numerical solution for different lattice sizes is shown in the upper panel of Fig. 6. Note that in this case the associated repeated quasienergy is not zero. Instead, they trace interesting curves in the complex plane, see the examples shown in the bottom panel of Fig. 6.
@@ V Bulk-boundary correspondence
For Hermitian systems, the bulk-boundary correspondence connects a topological invariant calculated for a system with periodic boundary conditions with the number of topologically protected edge modes in a system with open boundaries which have energies which converge exponentially to zero with system size. It is well-known that this type of bulk-boundary correspondence breaks down in the non-Hermitian case [16, 14, 21]. Some authors have tried to devise new topological invariants based on, for example, ’non-Bloch theory’ [14] but these attempts lack a solid mathematical foundation. Furthermore, these theories are based on translational invariance, however, ”topological features must be properties of a translationally not invariant Hamiltonian” [56] and thus should be stable against perturbations which break translational invariance. From a mathematical point of view, none of this is at all surprising: it is well-known from Toeplitz theory that topological invariants are in general related to properties of the singular value, not the eigenvalue, spectrum. The importance of the singular value spectrum in the non-Hermitian case has been pointed out, for example, already in Refs. [37, 36] but without establishing a rigorous bulk-boundary correspondence. An analysis of the singular value spectrum has also turned out to be useful in the classification of non-Hermitian random matrices [57]. It is also worth noting that another way of understanding the importance of singular values is Hermitization [58] where the hermitian block matrix
$\bar{H}=\begin{pmatrix}0&H\\ H^{\dagger}&0\end{pmatrix}$
is constructed. A point gap in $H$ then becomes a spectral gap in $\bar{H}$ and the positive eigenvalues of $\bar{H}$ correspond to the singular values of $H$. However, this approach does introduce an unphysical additional chiral symmetry and is not needed here. The case where the Hamiltonian $H$ itself is Hermitian is special because the singular values are then just the absolute values of the eigenvalues so that index theorems for singular values also apply to the eigenvalues. Another important and so far overlooked issue is that the eigenvalues of a finite non-Hermitian matrix do, in general, not converge to the eigenvalues of the semi-infinite case. This is similar to a discontinuous point.
All of these aspects can be fully understood if one starts from known index theorems in Toeplitz theory. In particular, the K-splitting theorem states [59] that a system with winding number $I$ has $K\geq|I|$ singular values which will converge to zero with increasing system size and which belong to states which are exponentially localized at the boundary. These topologically protected singular values are separated by a gap from the rest of the singular value spectrum. In the semi-infinite case, these states become exact eigenstates with exactly zero energy but for finite system size they only get mapped exponentially close to zero by the Hamiltonian but are not exact eigenstates. I.e., in a finite non-Hermitian system the topological zero-energy edge modes are ’hidden’, in general, and instead constitute very long-lived metastable states. The latter aspect is extremely important for a proper understanding of the topological properties of non-Hermitian systems and for testing these properties experimentally. One possible scenario for such an experiment is to prepare a system in a hidden zero mode—which is localized at one of the boundaries—in a cold atomic gas, trapped ion system, or some other quantum simulation/computation platform and to monitor a generalized Loschmidt echo. This Loschmidt echo will stay close to $1$ for a time which is proportional to system size thus indicating the metastable character and its convergence to a true eigenstate in the thermodynamic limit. The corresponding protocol is discussed in more detail in the Suppl. Mat. of Ref. [22]. We note that the required postselection to trajectories without jumps has already been performed for a single qubit [6]. Such an experiment, generalized to multiple qubits, would be able to detect the hidden zero modes.
We stress again that the formalism developed in Ref. [22] does encompass the standard bulk-boundary correspondence in the Hermitian case because for a Hermitian matrix the singular values are just the absolute values of the eigenvalues. Here we want to explicitly and fully analytically demonstrate this general bulk-boundary correspondence for the first time for the physically relevant minimal non-Hermitian model under consideration.
@@ V.1 Singular values
with boundaries fixed by the last line in (V.1). Using the method reviewed in App. B, we find that there are two solutions to the difference equations (V.1).
@@
The spectrum of (V.1) is given by the union of $\sigma_{\phi}^{2}$ and $\sigma_{\omega}^{2}$. Remarkably, the pairs of variables $\{V_{L},W_{R}\}$ and $\{V_{R},W_{L}\}$ decouple from each other showing the value of an analytical solution. This means that the characteristic polynomial associated with (V.1) factorizes into two independent factors depending solely on each of these pairs.
The transcendental equations (V.7) and (V.11) can be solved numerically. We find that (V.7) has $N/2$ real solutions in the interval $0<\phi<\pi$ when $|V_{L}/W_{R}|\geq 1$. When $|V_{L}/W_{R}|<1$ a complex conjugate pair emerges near $\pi$. This complex root produces an exponentially small singular value $\sigma_{\phi}$. Similarly, we find that (V.11) has $N/2$ real solutions in the interval $0<\omega<\pi$ when $|V_{R}/W_{L}|\geq 1$ and N/21N/2-1 real solutions plus a complex conjugate pair solution when $|V_{R}/W_{L}|<1$. In Fig. 7, we plot the smallest two singular values coming from each transcendental equation considering a lattice with $N=100$ sites.
@@
The Bloch Hamiltonian for the corresponding periodic problem with $u=0$ has block structure, see Eq. (III.30). There are therefore two winding numbers $\nu_{1,2}$ defined in Eq. (III.31). The K-splitting theorem then predicts that there are $K\geq|\nu_{1}|+|\nu_{2}|$ singular values $\sigma$ with $\lim_{N\to\infty}\sigma=0$ [59, 22]. This is fully consistent with our results; compare Fig. 7 with the phase diagram in the left panel of Fig. 2.
@@ V.2 Topological protection and hidden zero modes
In the limit of a semi-infinite chain, the system will have $K\geq|\nu_{1}|+|\nu_{2}|$ topologically protected edge modes with exactly zero energy. However, these states will only be eigenstates in this limit but not for a finite system. For a finite system, the singular values have to be considered as we have done above. Here we want to show explicitly how the exact edge states in the semi-infinite limit arise and how they become ’hidden’ if the system is finite.
@@
The condition $Hv=0$ leads to $v_{2}=0$ and $W_{R}v_{2j}+V_{L}v_{2j+2}=0$ implying that all even vector components are zero, $v_{2j}=0$. For the odd vector components we find the recurrence relation $V_{R}v_{2j-1}+W_{L}v_{2j+1}=0$ with the solution $v_{2j+1}=\left(-\frac{V_{R}}{W_{L}}\right)^{j}v_{1}$. In addition, we have to demand that the solution is normalizable, $v||^{2}=|v_{1}|^{2}\sum_{j=0}^{\infty}\left(-\frac{V_{R}}{W_{L}}\right)^{2j}=1$. This is possible if $|W_{L}|>|V_{R}|$. In this case, the semi-infinite chain has a zero energy edge mode $v$ exponentially localized at the left edge with vector components
@@
Again, for a finite system the only solution to $\tilde{H}w=0$ is $w=0$. The non-trivial edge mode with energy zero only exists in the thermodynamic limit. We note that we find the exact same separation of parameters as for the singular values: the first zero mode depends only on $\{V_{R},W_{L}\}$ and the second zero mode only on $\{V_{L},W_{R}\}$. For a finite system, we can truncate the vectors $v,w$. They are then no longer eigenvectors, $Hv\neq 0$, but they do get mapped exponentially close to zero, i.e. $Hv||\to 0$ for $N\to\infty$ and similarly for $w$. This is what we mean by hidden zero modes.
For the example considered in Fig. 8 with $V_{R}=W_{R}=1$ this means that the first zero mode is stable for $1<V_{L}<1$ and the second zero mode for $|W_{L}|>1$. This is fully consistent with the singular values as well as with the winding numbers as is expected based on the K-splitting theorem. We thus do have explicitly confirmed the bulk-boundary correspondence in this case.
To check that the singular values with $\lim_{N\to\infty}\sigma=0$ are topologically protected while zero eigenvalues are, in general, not protected for a finite non-Hermitian system, we also show in Fig. 8 both quantities when calculated for a finite Hamiltonian $H$ ($N=200$) with a random complex matrix $A$ with elements $|a_{ij}|\leq 0.02$ added. The results show that the singular values are stable and thus indeed topologically protected while the eigenvalues are, in general, not protected. More specifically, we find that for the case $W_{L}=0.5$ shown in the lower left panel there are two degenerate zero eigenvalues in the unperturbed system for $0.4\lesssim V_{L}\lesssim 0.5$. However, the small perturbation immediately moves them away from zero energy. They are not protected.
The situation is slightly more complicated in the case $W_{L}=1.5$ shown in the lower right panel of Fig. 8. In the unperturbed case, there are two degenerate zero energy eigenvalues for $1.3\lesssim V_{L}\lesssim 1.5$. Interestingly, adding a perturbation does not completely remove them from zero energy. Instead, they are stable for $1<V_{L}<1$. This can be understood as follows: in this regime, the non-Hermitian Hamiltonian with $V_{R}=W_{R}=1$ and $W_{L}=1.5$ can be adiabatically connected to a chiral Hermitian Hamiltonian without closing the gap. This can be achieved, for example, by $W_{R}\to 1.5$ and then $V_{R}\to V_{L}$. For the obtained Hermitian Hamiltonian one can calculate the winding number of the upper block which turns out to be $I=1$. The standard bulk-boundary correspondence then predicts the existence of two protected zero-energy edge modes. Since the gap never closes and the sub-lattice symmetry remains intact, these zero modes survive in the non-Hermitian case. However, the full topology of the non-Hermitian model also in this case is only captured by the singular values, not by the eigenvalues.